A rectangular stage has an area of 600 square yards. The length of the stage must be more than 10 yard more than the width. Find the dimension of the stage nedded.
Answer by usyim88hk(158)

Let x be the width of the stage:
x*(x+10) = 600
(the width times the length whcih is equal to 600)
x^2+10x = 600
x^2+10x-600 = 0
Now factor the equation
(x+30)(x-20) = 0
x = -30 or 20
Because length of anything can not be a negative number, so 20 is the width and x+10 whcih is 30 is the length of the stage.